Mr. Jones has $0.70 in dimes and nickels. The total number of coins is 2 more than 3 times the number of nickels. How many coins of each type does he have?
Answer by ikleyn(52863)

.
The condition literally means that the number of dimes is TWO times the number of nickels PLUs 2:


    D = 2N + 2.


Then the money equations is


    5N + 10*(2N + 2) = 70   cents

    5N + 20N + 20 = 70

    25N = 70-20 = 50

    N = 50/25 = 2.


Answer.  2 nickels and  2*2+2 = 6 dimes.

You may check it on your own.

-----------------

On coin problems,  see the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.

===============

Your problem has an original, fresh and unexpected formulation, and I like it - Thanks for posting it !